OpenStudy (anonymous):

Claire lives 6/8 of a mile from the store and 4/5 of this distance from her best friend. How far does Claire live from her best friend?

OpenStudy (anonymous):

of = multiplication

OpenStudy (anonymous):

What this is saying is that Claire's best friend lives 4/5 of the distance from Claire to the store. Since the distance from Claire's house to the store is 6/8 of a mile, the distance between Claire and her friend is just \[\frac{4}{5}\times \frac{6}{8}=\frac{24}{40}\]

OpenStudy (anonymous):

which you then simplify (like I showed you earlier).

OpenStudy (anonymous):

The answer should be 3/5 of a mile.

OpenStudy (anonymous):

thank you again

OpenStudy (anonymous):

You're welcome.